Grade 12th passMagical Mathematics[Interesting Approach]

If x,y,z>0 are in H.P., then what is the value of (x-y)/(y-z)?

Koyel Das
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student
if x, y, z are in H.P.
1/x, 1/y, 1/z are in A.P.
1/y –1/x = 1/z –1/y
(x – y) / xy = (y – z) / yz
hence
(x –y) / (y –z) = xy / yz = x /z
Regards
Arun (askIITians forum expert)